2,6,12,20.........
Answer by Cromlix(4381)

2,6,12,20,30,42,56
The difference between ascends
in units of two starting with 4
2-6 = 4
6-12 = 6
12-20 = 8
20-30 = 10
30-42 = 12
And so on....
Hope this helps.
:-)